A spring 20cm long is stretched to 25cm by a load of 50N. What will be it's length when stretched by 100N assuming that the elastic limit is not reached



1
Expert's answer
2022-03-03T13:23:18-0500

The spring constant can be determined from the Hook's law:



"k = \dfrac{F_1}{\Delta l_1} = \dfrac{50N}{25cm - 20cm} =10N/cm"

Then, the expansion under the second force:



"\Delta l_2 = \dfrac{F_2}{k} = \dfrac{100N}{10N/cm} =10cm"

Thus, the length of the spring will be:


"l_2 = 20cm+10cm = 30cm"

Answer. 30 cm.